# Improving Parallel Program Performance with LLM Optimizers via Agent-System Interface

### v2

@@ Abstract
Modern scientific discovery increasingly relies on high-performance computing for complex modeling and simulation. A key challenge in improving parallel program performance is efficiently mapping tasks to processors and data to memory, a process dictated by intricate, low-level system code known as mappers. Developing high-performance mappers demands days of manual tuning, posing a significant barrier for domain scientists without systems expertise. We introduce a framework that automates mapper development with generative optimization, leveraging richer feedback beyond scalar performance metrics. Our approach features the Agent-System Interface, which includes a Domain-Specific Language (DSL) to abstract away low-level complexity of system code and define a structured search space, as well as AutoGuide, a mechanism that interprets raw execution output into actionable feedback. Unlike traditional reinforcement learning methods such as OpenTuner, which rely solely on scalar feedback, our method finds superior mappers in far fewer iterations. With just 10 iterations, it outperforms OpenTuner even after 1000 iterations, achieving 3.8×3.8$\times$3.8 × faster performance. Our approach finds mappers that surpass expert-written mappers by up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks while reducing tuning time from days to minutes.
@@ 1 Introduction
For the second challenge, we adopt the generative optimization approach, a recent advance in optimization techniques. Unlike traditional methods such as reinforcement learning (Ansel et al., 2014), which rely solely on scalar rewards, generative optimization can utilize richer forms of feedback, such as error explanations and actionable suggestions expressed in natural language. This agentic optimization workflow has previously proven to be effective across various domains (Nie et al., 2024; Cheng et al., 2024; Yang et al., ; Khattab et al., 2023; Yuksekgonul et al., 2024). Our work is the first to apply such technique to the domain of system optimization.
@@ 3 Problem Definition
Motivation and Challenges
Search Space and Performance Impact
4 Our Approach: Agent-System Interface
@@ 4.1 Domain-Specific Language Design
As illustrated in Figure 2, the complexity of DSL code is significantly lower than that of C++. Figure 2(a) provides an example of a DSL mapper, highlighting the key features of our DSL. In contrast, Figure 2(b) shows a snippet from a C++ mapper, emphasizing the intricacy of low-level implementation details. According to Table A1, using the DSL results in an average lines of code reduction of 14×14$\times$14 ×. This substantial reduction makes DSL a more suitable target for LLM code generation, as it abstracts away the complexities inherent in low-level systems. As we will show in Section 5.2, LLMs generate DSL code more effectively, despite DSL having no examples in LLM training corpora, whereas C++ is widely represented.
@@ 4.2 Generative Optimization via AutoGuide
Figure 3: Agent Optimization Process. The mapper agent takes server specifications and application-specific information as input, generates mapper code, and executes it alongside the application on the server. Raw execution feedback is enriched using the AutoGuide mechanism and iteratively refined by an LLM optimizer to improve performance.
Case Raw Execution Output AutoGuide
Explain Suggest
@@
We present the optimization process in Figure 3. The agent takes two inputs: server specifications (e.g., CPU/GPU counts) and application information (e.g., task lists, data arguments). It generates mapper code, which is executed alongside the application code on the server. Raw execution feedback from the runtime is augmented with the AutoGuide mechanism and fed back to the LLM, iteratively refining the agent for improved mapper code generation.
@@ 5.1 Speedup of Application Performance
We evaluate our approach using 9 benchmarks. Circuit (Bauer et al., 2012) is a simulation benchmark that models electrical circuit behavior by simulating currents and voltages across interconnected nodes and wires. Stencil (Van der Wijngaart & Mattson, 2014) simulates a 2D grid where each point’s value is updated based on a stencil pattern determined by its neighbors. Pennant (Ferenbaugh, 2015) models unstructured mesh Lagrangian staggered-grid hydrodynamics, commonly used for simulating compressible flow. The remaining six benchmarks – Cannon’s (Cannon, 1969), SUMMA (Van De Geijn & Watts, 1997), PUMMA (Choi et al., 1994), Johnson’s (Agarwal et al., 1995), Solomonik’s (Solomonik & Demmel, 2011), and COSMA (Kwasniewski et al., 2019) – are widely studied parallel matrix multiplication algorithms, which we discuss in more detail in Section A.3.
@@
Agent-Optimized Mappers. Using Trace (Cheng et al., 2024), we evaluated the Trace and OPRO (Yang et al., ) search algorithms, running 10 iterations per application. To account for stochastic output, we repeated the process 5 times and report the average. The best mapper from Trace across runs is also reported.
@@ Results
We use normalized throughput as the performance metric in Figure 4, where higher values indicate better performance. The throughput is normalized relative to the expert-written mappers. All the best mappers found by Trace can match or surpass the expert-written mappers, underscoring the effectiveness of agent-based generative optimizer. Random mappers consistently exhibit low performance across all applications, emphasizing the critical role of mapping decisions. When comparing optimization trajectories, Trace performs similarly to OPRO, and significantly outperforms OpenTuner.
To further compare the agent-based optimizer with traditional reinforcement learning, we extended OpenTuner’s optimization iterations from 10 to 1000, as shown in Figure 5, where the x-axis is the log-scale of iterations and the y-axis represents relative throughput (averaged across all 9 benchmarks). Notably, Trace achieves a 3.8×3.8$\times$3.8 × speedup over OpenTuner even when OpenTuner is run for 1000 iterations. When both are limited to 10 iterations, Trace outperforms OpenTuner by $\times1$, demonstrating its ability to quickly identify high-performance mappings. This highlights the superiority of Trace (generative optimizer) over OpenTuner (traditional reinforcement learning). Moreover, Trace completes the entire optimization process in just 10 minutes per application, reducing mapper development time from days to minutes.
Case Analysis
@@
5.2 Ablation Study of DSL for Code Generation
In Section 5.1, we demonstrate the overall effectiveness of our approach. Here, we conduct an ablation study on the DSL, the core of the Agent-System Interface. Since successful generation is the foundation of optimization, this subsection focuses on how well the DSL helps LLMs generate correct mappers compared to C++, rather than directly optimizing performance.
@@
Table 2 shows that DSL achieves significantly higher generation success rates than C++ in both the single-trial and iterative refinement settings. This demonstrates the effectiveness of DSL’s design in abstracting system complexity and providing a high-level interface that enables LLMs to tackle complex system challenges in code generation. Incorporating iterative refinement with compiler feedback further improves success rates, resolving three compilation errors in C++ and two in the DSL. However, the gap between DSL mappers and C++ mappers remains substantial. Notably, these results are striking given that the DSL is a low-resource language with no pre-training or fine-tuning data, while C++ code is widely present in LLM training corpora.
@@
LLMs perform better with the DSL for two reasons. First, the semantic gap between natural language and code is smaller with the DSL than with C++. For example, writing a mapper to “align all data to 64 bytes in memory and use Fortran ordering” requires one line Layout * * * Align==64 F_order in the DSL because of its declarative design. In contrast, the C++ mapping API requires a sequence of operations to enforce alignment and ordering, which widens the semantic gap. Second, the DSL reduces the amount of code. As shown in Table A1, LLMs achieve an average reduction of 14×14$\times$14 × in lines of code, simplifying code generation. These results underscore the importance of a high-level agent-system interface.
@@ 5.3 Ablation Study of the AutoGuide Feedback
Figure 6 demonstrates that the full feedback mechanism consistently outperforms all reduced feedback variants. The 0-shot and 5-shot results perform the worst, underscoring the importance of feedback-based iterative refinement. This highlights the value of an agentic workflow, showing that performance improvements are not solely driven by prompting the LLM but are a direct result of the iterative refinement inherent in the workflow design.
@@ Acknowledgements
We thank Yuhui Zhang, Chenyang Yang, Genghan Zhang, Qizheng Zhang, Mohammad R. Fadiheh, Ed Chen, Mert Yuksekgonul, Jiaxin Ge, Xinyan Hu, Tingyu Zhu, Simon Guo, Anne Ouyang, and Adhish Ramkumar for the feedback. We thank Ching-An Cheng and Ken Liu for their discussions and feedback.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Illustration for Mapping
A.2 Tables for Lines of Code Comparison
We show the lines of code comparison between DSL and C++ mappers in Table A1.
Application 1 2 3 4 5 6 7 8 9 Avg.
LoC in C++ 347 306 379 447 437 430 428 433 448 406
LoC in DSL 16 14 16 38 38 38 33 38 32 29
LoC Reduction $\times2$ $\times2$ 24×24$\times$24 × 12×12$\times$12 × 12×12$\times$12 × $\times1$ 13×13$\times$13 × $\times1$ 14×14$\times$14 × 14×14$\times$14 ×
Table A1: Lines of Code (LoC) comparison between DSL and C++ mappers. The DSL achieves a 14×14$\times$14 × average reduction in LoC, making it a more suitable target for LLM code generation.
@@ A.3 Parallel Matrix Multiplication Algorithms
A.4 Examples of Feedback Configurations
@@
A.5 Trace Agent Code
@@ A.6 Mapping Strategies
A.7 Generated Mapper Examples
@@
7def block1d(Task task) {
8 ip = task.ipoint;
9 return mgpu[ip[0] % mgpu.size[0], ip[0] % mgpu.size[1]];
10}
11
12IndexTaskMap task_1 block1d;
13
14def cyclic1d(Task task) {
15 ip = task.ipoint;
16 linearize = ip[0] * 2 + ip[1];
17 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
18}
19
20IndexTaskMap task_1 cyclic1d;
21
22def cyclic2d(Task task) {
23 ip = task.ipoint;
24 linearize = ip[0] + ip[1] * 2;
25 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
26}
27
28IndexTaskMap task_1 cyclic2d;
29
30def linearize3D(Task task) {
@@
34}
35
36IndexTaskMap task_1 linearize3D;
37
38def linearize2D(Task task) {
@@
42}
43
44IndexTaskMap task_1 linearize2D;
45IndexTaskMap task_2 block1d;
46IndexTaskMap task_2 cyclic1d;
47IndexTaskMap task_2 cyclic2d;
48IndexTaskMap task_2 linearize3D;
49IndexTaskMap task_2 linearize2D;
50IndexTaskMap task_3 block1d;
51IndexTaskMap task_3 cyclic1d;
52IndexTaskMap task_3 cyclic2d;
53IndexTaskMap task_3 linearize3D;
54IndexTaskMap task_3 linearize2D;
55IndexTaskMap task_5 block1d;
56IndexTaskMap task_5 cyclic1d;
57IndexTaskMap task_5 cyclic2d;
58IndexTaskMap task_5 linearize3D;
59IndexTaskMap task_5 linearize2D;
60
61m_2d = Machine(GPU);
62def same_point(Task task) {

### v4

@@ Abstract
Modern scientific discovery increasingly relies on high-performance computing for complex modeling and simulation. A key challenge in improving parallel program performance is efficiently mapping tasks to processors and data to memory, a process dictated by intricate, low-level system code known as mappers. Developing high-performance mappers demands days of manual tuning, posing a significant barrier for domain scientists without systems expertise. We introduce a framework that automates mapper development with generative optimization, leveraging richer feedback beyond scalar performance metrics. Our approach features the Agent-System Interface, which includes a Domain-Specific Language (DSL) to abstract away the low-level complexity of system code and define a structured search space, as well as AutoGuide, a mechanism that interprets raw execution output into actionable feedback. Unlike traditional reinforcement learning methods such as OpenTuner, which rely solely on scalar feedback, our method finds superior mappers in far fewer iterations. With just 10 iterations, it outperforms OpenTuner even after 1000 iterations, achieving 3.8×3.8$\times$3.8 × faster performance. Our approach finds mappers that surpass expert-written mappers by up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks while reducing tuning time from days to minutes.
@@ 1 Introduction
For the second challenge, we adopt the generative optimization approach, a recent advance in optimization techniques. Unlike traditional methods such as reinforcement learning (Ansel et al., 2014), which rely solely on scalar rewards, generative optimization can utilize richer forms of feedback, such as error explanations and actionable suggestions expressed in natural language. This agentic optimization workflow has previously proven to be effective across various domains (Nie et al., 2024; Cheng et al., 2024; Yang et al., 2023; Khattab et al., 2023; Yuksekgonul et al., 2024). Our work is the first to apply such technique to the domain of system optimization.
@@ 14 TaskSlice slice;
Motivation and Challenges
Search Space and Performance Impact
4 Our Approach: Agent-System Interface
@@ 4.1 Domain-Specific Language Design
Figure 3: Agent Optimization Process. The mapper agent takes server specifications and application-specific information as input, generates mapper code, and executes it alongside the application on the server. Raw execution feedback is enriched using the AutoGuide mechanism, and the mapper is iteratively refined by an LLM optimizer to improve performance.
Case Raw Execution Output AutoGuide
Explain Suggest
As illustrated in Figure 2, the complexity of DSL code is significantly lower than that of C++. Figure 2(a) provides an example of a DSL mapper, highlighting the key features of our DSL. In contrast, Figure 2(b) shows a snippet from a C++ mapper, emphasizing the intricacy of low-level implementation details. Across the benchmarks, using the DSL results in an average lines of code reduction of 14×14$\times$14 ×. This substantial reduction makes DSL a more suitable target for LLM code generation, as it abstracts away the complexities inherent in low-level systems. As we will show in Section 5.2, LLMs generate DSL code more effectively, despite DSL having no examples in LLM training corpora, whereas C++ is widely represented.
@@
Our DSL is designed to express a wide range of high-performance mapping strategies, including all of the most important decisions. While there may be cases where certain optimizations are not directly expressible, we have not encountered any. Despite being more constrained than general-purpose C++, the DSL has been proven to be effective: all mappers discovered by our agent that outperform expert-written C++ implementations are expressible within the current DSL.
@@ 4.2 Generative Optimization via AutoGuide
We present the optimization process in Figure 3. The agent takes two inputs: server specifications and application metadata. Server specifications detail the hardware configuration, including the number of CPUs and GPUs per node, as well as the total node count. Application metadata provides information on task names and the associated data arguments accessed by each task. These inputs define the structured search space explored by the agent during optimization. The agent, using the given inputs, generates mapper code that is executed alongside the application code on the server. Raw execution feedback from the runtime is augmented with the AutoGuide mechanism and fed back to the LLM, iteratively refining the agent for improved mapper code generation.
@@ 5.1 Speedup of Application Performance
Benchmarks
Our evaluation utilizes a suite of 9 benchmarks, including 3 scientific computing workloads and 6 well-known matrix multiplication algorithms. Circuit is a simulation benchmark that models electrical circuit behavior by simulating currents and voltages across interconnected nodes and wires (Bauer et al., 2012). Stencil simulates a 2D grid where each point’s value is updated based on a stencil pattern determined by its neighbors (Van der Wijngaart & Mattson, 2014). Pennant models unstructured mesh Lagrangian staggered-grid hydrodynamics, commonly used for simulating compressible flow (Ferenbaugh, 2015). The remaining six benchmarks – Cannon’s, SUMMA, PUMMA, Johnson’s, Solomonik’s, and COSMA – are well-known parallel matrix multiplication algorithms (Cannon, 1969; Van De Geijn & Watts, 1997; Choi et al., 1994; Agarwal et al., 1995; Solomonik & Demmel, 2011; Kwasniewski et al., 2019). Parallel matrix multiplication remains an active research topic due to its central role in high-performance computing and scientific simulations (Yadav et al., 2022). Furthermore, improving matrix multiplication performance has a broad impact, as it accelerates numerous downstream machine learning workloads (Jangda et al., 2022; Zheng et al., 2025). We discuss these matrix multiplication algorithms in more detail in Section A.3. This benchmark suite provides both depth with its representative matrix multiplication algorithms and variety with its range of scientific computing workloads.
@@
Agent-Optimized Mappers. Using Trace (Cheng et al., 2024), we evaluated the Trace and OPRO (Yang et al., 2023) search algorithms, running 10 iterations per application. To account for stochastic output, we repeated the process 5 times and report the average. The best mapper from Trace across runs is also reported.
@@ Results
In Figure 4, we use normalized throughput as our performance metric, where higher values indicate better performance. The throughput is normalized relative to the expert-written mappers, providing a clear baseline for comparison. Our focus is on measuring end-to-end performance, which includes both the correctness and efficiency of the generated mappers. If the generated code has any syntax or runtime issues, its throughput is recorded as 0. We report the best mappers found by Trace, and the average optimization trajectories of Trace, OPRO and OpenTuner over 10 iterations across 5 runs.
All the best mappers found by Trace can match or surpass the expert-written mappers, underscoring the effectiveness of agent-based generative optimizer. In our context, reporting the best-performing mapper is appropriate. Mapper optimization is an offline process, and in practice, it is standard to run the optimizer multiple times and deploy the best result. Once identified, the mapper can be reused across repeated executions on the same application, input, and hardware, incurring no further search cost.
@@
Random mappers consistently exhibit low performance across all applications, emphasizing the critical role of mapping decisions. For each application, we generate 10 random mappers by sampling from the full DSL-defined search space, totaling 90 mappers across 9 applications. Among them, 74 (82.2%) raise runtime errors due to invalid mapping decisions. The runtime system enforces correctness by rejecting such mappers during execution, resulting in a throughput of 0.
When comparing optimization trajectories, Trace performs similarly to OPRO, and significantly outperforms OpenTuner. To further compare the agent-based optimizer with traditional reinforcement learning, we extended OpenTuner’s optimization iterations from 10 to 1000, as shown in Figure 5, where the x-axis is the log-scale of iterations and the y-axis represents relative throughput (averaged across all 9 benchmarks). Notably, Trace achieves a 3.8×3.8$\times$3.8 × speedup over OpenTuner even when OpenTuner is run for 1000 iterations. When both are limited to 10 iterations, Trace outperforms OpenTuner by $\times1$, demonstrating its ability to quickly identify high-performance mappings. This highlights the superiority of Trace (generative optimizer) over OpenTuner (traditional reinforcement learning). Moreover, Trace completes the entire optimization process in just 10 minutes per application, reducing mapper development time from days to minutes.
To offer a more comprehensive view of performance variations, we present additional statistics, including the mean, standard deviation, worst, median, and best normalized throughput for both our method and OpenTuner. These statistics are derived from five runs for each benchmark, as detailed in Section A.4.
Case Analysis
5.2 Ablation Study of DSL for Code Generation
In Section 5.1, we demonstrate the overall effectiveness of our approach. Here, we conduct an ablation study on the DSL, the core of the Agent-System Interface. Since successful generation is the foundation of optimization, this subsection focuses on how well the DSL helps LLMs generate syntactically and semantically correct mappers compared to C++, rather than directly optimizing performance.
@@
Table 2 shows that DSL achieves significantly higher generation success rates than C++ in both the single-trial and iterative refinement settings. This demonstrates the effectiveness of DSL’s design in abstracting system complexity and providing a high-level interface that enables LLMs to tackle complex system challenges in code generation. Incorporating iterative refinement with compiler feedback further improves success rates, resolving four compilation errors in C++ and two in the DSL. However, the gap between DSL mappers and C++ mappers remains substantial. Notably, these results are striking given that the DSL is a low-resource language with no pre-training or fine-tuning data, while C++ code is widely present in LLM training corpora.
@@
LLMs perform better with the DSL for two reasons. First, the semantic gap between natural language and code is smaller with the DSL than with C++. For example, writing a mapper to “align all data to 64 bytes in memory and use Fortran ordering” requires one line Layout * * * Align==64 F_order in the DSL because of its declarative design. In contrast, the C++ mapping API requires a sequence of operations to enforce alignment and ordering, which widens the semantic gap. Second, the DSL reduces the amount of code. As discussed before, LLMs achieve an average reduction of 14×14$\times$14 × in lines of code, simplifying code generation. These results underscore the importance of a high-level agent-system interface.
@@ 5.3 Ablation Study of the AutoGuide Feedback
Figure 6 demonstrates that the full feedback mechanism consistently outperforms all reduced feedback variants. The 0-shot and 5-shot results perform the worst, underscoring the importance of feedback-based iterative refinement. This highlights the value of an agentic workflow, showing that performance improvements are not solely driven by prompting the LLM but are a direct result of the iterative refinement in the workflow design.
@@ Acknowledgements
This work was supported in part by a Google Research Award. We thank Yuhui Zhang, Genghan Zhang, Qizheng Zhang, Mohammad R. Fadiheh, Ed Chen, Mert Yuksekgonul, Chenyang Yang, Jiaxin Ge, Simon Guo, and Anne Ouyang for the feedback. We thank Ching-An Cheng and Ken Liu for their discussions and feedback.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Illustration for Mapping
A.2 DSL Grammar
Terminals: TaskName, RegionName, var, int
Grammar Rules:
[…]
@@ A.3 Parallel Matrix Multiplication Algorithms
A.4 Additional Performance Statistics
In our setting, reporting the best result across multiple runs is appropriate, as the best mapper is the one that is desired by the user. Mapper search is an offline optimization process, and it is feasible to run the optimizer multiple times to select the highest-performing mapper. Once identified, this mapper can be reused without incurring additional search cost, as the deployment scenario (application, input, and hardware) remains fixed.
That said, additional statistics on performance varations can provide a more complete picture. Here we include the mean, standard deviation, worst, median, and best normalized throughput across five runs for each benchmark of our approach Trace and OpenTuner.
Benchmark Mean Std Dev Worst Median Best
Circuit 1.33$\times$ 0.01 1.31$\times$ 1.33$\times$ 1.34$\times$
Stencil 1.01$\times$ 0.01 1.00$\times$ 1.01$\times$ 1.02$\times$
Pennant 1.03$\times$ 0.02 1.00$\times$ 1.03$\times$ 1.04$\times$
Cannon 1.09$\times$ 0.00 1.08$\times$ 1.09$\times$ 1.09$\times$
SUMMA 0.86$\times$ 0.48 0.00$\times$ 1.07$\times$ 1.09$\times$
PUMMA 0.57$\times$ 0.55 0.00$\times$ 0.66$\times$ 1.09$\times$
Johnson 0.98$\times$ 0.17 0.68$\times$ 1.06$\times$ 1.07$\times$
Solomonik 0.52$\times$ 0.41 0.00$\times$ 0.61$\times$ 1.09$\times$
COSMA 1.25$\times$ 0.03 1.23$\times$ 1.23$\times$ 1.31$\times$
Table A1: Normalized throughput of our framework Trace.
Benchmark Mean Std Dev Worst Median Best
Circuit 0.97$\times$ 0.16 0.81$\times$ 0.99$\times$ 1.20$\times$
Stencil 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
Pennant 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
Cannon 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
SUMMA 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
PUMMA 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
Johnson 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
Solomonik 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
COSMA 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
Table A2: Normalized throughput of OpenTuner.
Our method achieves relatively stable performance across most benchmarks. The higher variance and occasional 0.00× worst-case throughput observed in SUMMA, PUMMA, and Solomonik are due to invalid mapper configurations in the search space (e.g., violating cuBLAS layout constraints). The runtime enforces correctness by rejecting such configurations during execution. While the generative optimizer typically learns to avoid these cases through the AutoGuide mechanism, occasional failures within the 10-iteration budget are still possible. In practice, such failures can be mitigated by repeating the optimization and selecting the best-performing mapper. In contrast, OpenTuner, despite running the same number of iterations, fails to generate valid mappers for 8 out of 9 benchmarks. This highlights the difficulty of exploring the search space using traditional reinforcement learning methods.
A.5 Examples of Feedback Configurations
@@
A.6 Trace Agent Code
@@ A.7 Mapping Strategies
A.8 Generated Mapper Examples
@@
7def linearize3D(Task task) {
@@
11}
12
13IndexTaskMap task_1 linearize3D;
14
15def linearize2D(Task task) {
@@
19}
20
21IndexTaskMap task_1 linearize2D;
22IndexTaskMap task_2 linearize2D;
23IndexTaskMap task_3 linearize2D;
24IndexTaskMap task_5 linearize2D;
25
26m_2d = Machine(GPU);
27def same_point(Task task) {